# Efficient Generation of Multi-partite Entanglement between Non-local Superconducting Qubits using Classical Feedback

### v1

@@ Appendix D Measurement-Induced Dephasing
Figure A3: Benchmarking the Impact of Mid-Circuit Measurements. To measure the impact of MCMs on idling spectator qubits, we perform CB on the spectator qubits with an interleaved layer consisting of a MCM on another qubit. Here, $B_{P}$ is rotation to +1 eigenstate of the basis of the Pauli $P$, $\mathcal{P}$ is a randomly sampled Pauli gate, $m$ is the number of interleaved cycles, and $B^{\dagger}_{C(P)}$ rotates the system back to the original basis state.
@@
Mid-circuit measurements (MCMs) can lead to unwanted dephasing on nearby idling spectator qubits. This can happen in two different ways: firstly, when an ancilla qubits is measured, the static ZZ coupling Mitchell et al. (2021) between coupled qubits can dephase the spectator qubit during the measurement process. In other words, because measurement dephases the ancilla qubit, the coupled data qubit similarly sees a $Z$-type error mediated via the ZZ coupling. This only results in dephasing between qubits that are directly coupled, and assumes that the ZZ coupling cannot be turned off during the MCM. This is the case for the qubits used in this experiment, as they are coupled via fix-frequency resonators in a ring geometry (see Fig. A1a). However in systems with tunable couplers, this ZZ coupling can be largely suppressed, which would help alleviate the dephasing seen by the spectator qubit.
@@
The other way in which MCMs can dephase spectator qubits is via unintentional direct measurement, i.e., measurement-induced dephasing Gambetta et al. (2006). When the readout resonator frequencies of nearby qubits are not well-resolved spectrally, the measurement of one qubit can inadvertently result in the measurement of other qubits. This is the case for multiple pairs of qubits on the current device (see Fig. A1c). For example, the readout resonators coupled to Q0 and Q1 are within $\sim$4 MHz of each other, and the readout resonators coupled to Q3, Q4, and Q5 are within $\sim$11 MHz of each other. This is not done intentionally; rather, it is the result of inaccuracies in the fabrication process. This directly limits which qubits can be used as ancillas for MCMs and which can be used as data qubits, because the measurement of the ancilla qubit can result in a complete loss of phase coherence on data qubit.
@@
To measure the impact of MCMs on neighboring spectator qubits, we utilize cycle benchmarking (CB) with a slight variation (see Fig. A3). Here, the interleaved cycle is a MCM on one qubit, with all of the other qubits left idling. However, because the goal is to quantify the impact of MCMs on spectator qubits, we only twirl on the spectator qubits, we do not twirl the qubit being measured. A similar strategy was done in Ref. Govia et al. (2023) using RB, but RB cannot distinguish the impact of errors along different measurement axes. The utility of CB is that it enables to quantify how a MCM affects other qubits along $X$, $Y$, and $Z$. Moreover, we can glean information from both the state-preparation and measurement (SPAM) constant $A_{P}$ for the exponential decay of each Pauli operator $P$ (termed “Pauli decays”), as well as from the exponential fit parameter $p_{P}$. If the effect of a MCM on an idle qubit is small, then one should be able to accurately fit $A_{P}$ and $p_{P}$. However, if the effect of a MCM is large enough such that no useable information can be gleaned for a given $P$, then $A_{P}$ will be small (close to zero) and $p_{P}$ cannot be trust (assuming an exponential can even be fit to the data, which is not always the case). This is the case if the MCM completely dephases the idle qubit, for which $A_{X}\approx 0$ and $A_{Y}\approx 0$ (see, e.g., the bottom row of Fig. A5). When an exponential can be fit to the data, $p_{P}$ gives the probability with which $P$ is preserved by the error process. More generally, any Pauli error $Q$ will reduce the probability $p_{P}$ for any non-commuting Pauli operator $P$. Thus, if a MCM causes a spectator qubit to dephase (which is a $Z$-type error), this will affect $p_{X}$ and $p_{Y}$ (but will not affect $p_{Z}$). Therefore, we often observe that $p_{Z}>p_{X},p_{Y}$, which is a strong indication of measurement-induced dephasing.
@@
Using CB, we systematically characterize the effect of MCMs on all nearest-neighbor idling qubits across our full device (see Figs. A4A7). We compare the effect of MCMs to the case when the spectator qubits idle (with no MCM) for the duration of the measurement, as well as the case when dynamically decoupling (DD; specifically, an XY8 sequence Maudsley (1986); Ahmed et al. (2013)) is applied to the spectator qubits during the MCM. In all cases, we observe that the MCM has at least some effect on the spectator qubits. If the effect is small, we observe that DD helps reduce the dephasing (i.e., increase $p_{X}$ and $p_{Y}$). However, if the effect is large, DD does not help and can even impact the Pauli $Z$ decay adversely (see, e.g., the bottom row of Fig. A5). This is because DD relies on the assumption that the dephasing noise is quasi-static (i.e., slowly varying) over the course of an experiment, such that one can “echo away” the phase error. However, when a MCM is strong enough to inadvertently measure a nearby qubit, this breaks the underlying assumption of DD, as measurement back action introduces strong stochastic phase rotations Kockum et al. (2012). In such cases, a DD sequence like XY8 will not return to the qubit to the original state, inevitably introducing $X$- and $Y$-type errors (and thus affecting the Pauli-$Z$ decays).
@@
Figure A4: Measurement-Induced Dephasing. We observe that the measurement of Q1 almost completely dephases Q0; this is expected, because the readout resonator frequencies of Q0 and Q1 are only separated by $\sim$4 MHz (see Fig. A1c). In this case, DD applied to Q0 has an adverse effect on the Pauli-$Z$ decay. However, the measurement of Q1 has a much smaller effect on Q2, and DD applied to Q2 helps mitigate the dephasing. In contrast, the measurement of Q2 has a small effect on both Q1 and Q3. DD applied to both qubits helps mitigate the dephasing.
@@
Figure A5: Measurement-Induced Dephasing. We observe that the measurement of Q3 has a small effect on Q2, but almost completely dephases Q4 and Q5 (note that the Pauli-$X$ and Pauli-$Y$ decays for Q4 could not even be properly fit). This is not surprisingly, because the readout frequencies of Q3, Q4, and Q5, are only separated by $\sim$11 MHz (see Fig. A1c). While DD does help with the dephasing on Q2, it has minimal impact on Q4 and Q5. Similarly, we observe that the measurement of Q4 strongly affects Q3, and completely dephases Q5. DD helps Q3, but completely decoheres Q5.
@@
Figure A6: Measurement-Induced Dephasing. Similar to what we observed in Fig. A5, the measurement of Q5 completely dephases Q4, and DD has an adverse effect. However, Q6 is only marginally affected by Q5, and DD does indeed help. However, on the contrary, the measurement of Q5 seems to mildly dephase Q5, and DD does not appear to help. This is surprising, because the qubit and readout frequencies spectrally well-resolved. The measurement of Q6 has a smaller effect on Q7, and DD only marginally helps.
@@
Figure A7: Measurement-Induced Dephasing. We observe that the measurement of Q7 has a small effect on Q7 and that DD helps mitigate the dephasing. Q7 has a much larger effect on Q6, but DD mitigates most of the effect. In contrast, the measurement of Q0 completely dephases Q1 and completely decoheres Q7. DD has an adverse effect on Q1 and no effect on Q7.

### v2

@@ Appendix E Measurement-Induced Dephasing
Figure 7: Benchmarking the Impact of Mid-Circuit Measurements. To measure the impact of MCMs on idling spectator qubits, we perform CB on the spectator qubits with an interleaved layer consisting of a MCM on another qubit. Here, $B_{P}$ is rotation to +1 eigenstate of the basis of the Pauli $P$, $\mathcal{P}$ is a randomly sampled Pauli gate, $m$ is the number of interleaved cycles, and $B^{\dagger}_{C(P)}$ rotates the system back to the original basis state.
@@
Mid-circuit measurements (MCMs) can lead to unwanted dephasing on nearby idling spectator qubits. This can happen in two different ways: firstly, when an ancilla qubits is measured, the static $ZZ$ coupling Mitchell et al. (2021) between coupled qubits can dephase the spectator qubit during the measurement process Jurcevic and Govia (2022) if the ancilla qubit undergoes a measurement-induced state transition Dumas et al. (2024). This only results in dephasing between qubits that are directly coupled, and assumes that the $ZZ$ coupling cannot be turned off during the MCM. This is the case for the qubits used in this experiment, as they are coupled via fix-frequency resonators in a ring geometry (see Fig. 5a). However in systems with tunable couplers, this $ZZ$ coupling can be largely suppressed, which would help alleviate the dephasing seen by the spectator qubit.
@@
The other way in which MCMs can dephase spectator qubits is via unintentional direct measurement, i.e., measurement-induced dephasing Gambetta et al. (2006). When the readout resonator frequencies of nearby qubits are not spectrally well-resolved, the measurement of one qubit can inadvertently result in the measurement of other qubits. This is the case for multiple pairs of qubits on the current device (see Fig. 5c). For example, the readout resonators coupled to Q0 and Q1 are within $\sim$4 MHz of each other, and the readout resonators coupled to Q3, Q4, and Q5 are within $\sim$11 MHz of each other. This is not done intentionally; rather, it is the result of inaccuracies in the fabrication process. This directly limits which qubits can be used as ancillas for MCMs and which can be used as data qubits, because the measurement of the ancilla qubit can result in a complete loss of phase coherence on data qubits.
@@
To measure the impact of MCMs on neighboring spectator qubits, we utilize cycle benchmarking (CB) with a slight variation (see Fig. 7). Here, the interleaved cycle is a MCM on one qubit, with all of the other qubits left idling. However, because the goal is to quantify the impact of MCMs on spectator qubits, we only twirl the spectator qubits, we do not twirl the qubit being measured. A similar strategy was done in Ref. 75 using RB, but RB cannot distinguish the impact of errors along different measurement axes. The utility of CB is that it enables one to quantify how a MCM affects other qubits along $X$, $Y$, and $Z$. Moreover, we can glean information from both the state-preparation and measurement (SPAM) constant $A_{P}$ for the exponential decay of each Pauli operator $P$ (termed “Pauli decays”), as well as from the exponential fit parameter $p_{P}$. If the effect of a MCM on an idle qubit is small, then one should be able to accurately fit $A_{P}$ and $p_{P}$. However, if the effect of a MCM is large enough such that no useable information can be gleaned for a given $P$, then $A_{P}$ will be small (close to zero) and $p_{P}$ cannot be trusted (assuming an exponential can even be fit to the data, which is not always the case). This is the case if the MCM completely dephases the idle qubit, for which $A_{X}\approx 0$ and $A_{Y}\approx 0$ (see, e.g., the bottom row of Fig. 9). When an exponential can be fit to the data, $p_{P}$ gives the probability with which $P$ is preserved by the error process. More generally, any Pauli error $Q$ will attenuate the probability $p_{P}$ for any non-commuting Pauli operator $P$ Hashim et al. (2024). Thus, if a MCM causes a spectator qubit to dephase (which is a $Z$-type error), this will affect $p_{X}$ and $p_{Y}$ (but will not affect $p_{Z}$). Therefore, we often observe that $p_{Z}>p_{X},p_{Y}$, which is a strong indication of measurement-induced dephasing.
@@
Using CB, we systematically characterize the effect of MCMs on all nearest-neighbor idling qubits across our full device (see Figs. 811). We compare the effect of MCMs to the case when the spectator qubits idle (with no MCM) for the duration of the measurement, as well as the case when dynamically decoupling (DD; specifically, an XY8 sequence Maudsley (1986); Ahmed et al. (2013)) is applied to the spectator qubits during the MCM. In all cases, we observe that the MCM has at least some effect on the spectator qubits. If the effect is small, we observe that DD helps reduce the dephasing (i.e., increase $p_{X}$ and $p_{Y}$). However, if the effect is large, DD does not help and can even impact the Pauli $Z$ decay adversely (see, e.g., the bottom row of Fig. 9). This is because DD relies on the assumption that the dephasing noise is quasi-static (i.e., slowly varying) over the course of an experiment, such that one can “echo away” the phase error. However, when a MCM is strong enough to inadvertently measure a nearby qubit, this breaks the underlying assumption of DD, as measurement back-action introduces strong stochastic phase rotations Kockum et al. (2012). In such cases, a DD sequence like XY8 will not return to the qubit to the original state, inevitably introducing $X$- and $Y$-type errors (and thus affecting the Pauli-$Z$ decays).
@@
Figure 8: Measurement-Induced Dephasing. We observe that the measurement of Q1 almost completely dephases Q0; this is expected, because the readout resonator frequencies of Q0 and Q1 are only separated by $\sim$4 MHz (see Fig. 5c). In this case, DD applied to Q0 has an adverse effect on the Pauli-$Z$ decay. However, the measurement of Q1 has a much smaller effect on Q2, and DD applied to Q2 helps mitigate the dephasing. In contrast, the measurement of Q2 has a small effect on both Q1 and Q3. DD applied to both qubits helps mitigate the dephasing.
@@
Figure 9: Measurement-Induced Dephasing. We observe that the measurement of Q3 has a small effect on Q2, but almost completely dephases Q4 and Q5 (note that the Pauli-$X$ and Pauli-$Y$ decays for Q4 could not even be properly fit). This is not surprising, because the readout frequencies of Q3, Q4, and Q5, are only separated by $\sim$11 MHz (see Fig. 5c). While DD does help with the dephasing on Q2, it has minimal impact on Q4 and Q5. Similarly, we observe that the measurement of Q4 strongly affects Q3, and completely dephases Q5. DD helps Q3, but completely decoheres Q5.
@@
Figure 10: Measurement-Induced Dephasing. Similar to what we observed in Fig. 9, the measurement of Q5 completely dephases Q4, and DD has an adverse effect. However, Q6 is only marginally affected by Q5, and DD does indeed help. However, on the contrary, the measurement of Q6 seems to mildly dephase Q5, and DD does not appear to help. This is surprising, because the qubit and readout frequencies spectrally well-resolved. The measurement of Q6 has a smaller effect on Q7, and DD only marginally helps.
@@
Figure 11: Measurement-Induced Dephasing. We observe that the measurement of Q7 has a small effect on Q0 and that DD helps mitigate the dephasing. Q7 has a much larger effect on Q6, but DD mitigates most of the effect. In contrast, the measurement of Q0 completely dephases Q1 and completely decoheres Q7. DD has an adverse effect on Q1 and no effect on Q7.